Elise is four times as old as Amanda. In ten years, Amanda's age will be half of Elise's. How old are they now?
Answer by usyim88hk(158)

I love this kind of question, let's start now
:
Let x be the age of Amanda, so now we know
:
x is Amanda's age (now)
4x is Elise's age (now)
x+10 is Amanda's age (after 10 years)
4x+10 or 2(x+10) is Elise's age (after 10 years)
:
So now we know that 4x+10 is equal to (x+10)/2 because they are both the age of Elise after ten years, so now we can make an equation and solve it
:
4x+10 = 2(x+10)
:
Divide by 2 on both sides
:
2x+5 = x+10
:
subtract x from both sides
:
x+5 = 10
:
Subtract 5 from both sides
:
x=5!!!
:
So now we know that Amanda is five years old now, and Elise is 20 years old now